OpenStudy (anonymous):

lx+6l less than or equal to 5

OpenStudy (maheshmeghwal9):

\[|x+6|\le 5 \implies \]\[x+6\le 5\]or\[x+6\ge -5.\]Now try to do:)

OpenStudy (anonymous):

It would be -1 and -11 I believe. x+6=5 x+6=-5

OpenStudy (maheshmeghwal9):

no u did wrong:/

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

x is less than or equal to -1 and x is greater than or equal to -11.

OpenStudy (maheshmeghwal9):

\[x \le -1.\]&\[x \ge-11.\]

OpenStudy (maheshmeghwal9):

u missed actually inequality signs that is very bad:/

OpenStudy (anonymous):

Write the solutions as either the union or the intersection of two sets